OpenStudy (anonymous):

solve sqrt (2x+3) - sqrt (x+1) = 1

OpenStudy (anonymous):

i got -3,-1

OpenStudy (anonymous):

-3 does not work

OpenStudy (anonymous):

so 3,-1

OpenStudy (anonymous):

if you squared when you were solving you have to check your answers, because you might have introduced an extraneous one while squaring

OpenStudy (anonymous):

chit, so 3

OpenStudy (anonymous):

3 works. did you check -1?

OpenStudy (anonymous):

i did, did not work

OpenStudy (anonymous):

oh it looks like it does work

OpenStudy (anonymous):

\[\sqrt{2(-1)+3}-\sqrt{-1+1}=\sqrt{1}-\sqrt{0}=1\]

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

yup -1 and 3 both work

OpenStudy (anonymous):

ty